An element x has atomic number of 29has two isotopes y And Z 34 and 38 respectively.. calculate the mass number for each isotopes and the nuclear symbol.. calculate the percentage abundance

1
Expert's answer
2021-11-16T11:23:02-0500

(34+29)+38-29

=63+9= 72/2= 36